# Pattern formation in a coupled driven diffusive system

### v1

@@ II.2 Results from simulations and comparisons with the lattice gas
To explore the behavior of this model, we carry out simulations on lattices of several $L$’s from $32$ to $256$. We vary the overall density, $\rho$, from $0.10$ to $0.95$. Here, we report only results from the extreme cases of $\delta=0$ and 11, which exhibit all the key phenomenology of the model at intermediate drive values. The main set of studies is initiated from a uniform initial state with $\rho_{A,B}\left(\mathbf{s}\right)=\rho/2$ and, for measurements in the steady-state, up to $100$ realizations per run, for up to $2\cdot 10^{7}$ MCS. With no drive ($\delta=0$), the behavior is essentially the same as the WRLG, displaying a disordered (phase-separated) state at low (high) densities, with the critical density estimated to be $\rho_{c}^{\left(0\right)}\simeq 0.39$. For $\rho$’s below this value, the structure factors are isotropic and display the standard Ornstein-Zernike form. Near criticality, we expect the properties to be in the Ising class [29] and did not pursue this issue. At densities above $\rho_{c}^{\left(0\right)}$, the usual coarsening dynamics is observed and we verified [48] that our system follows the Lifshitz-Slyozov law, i.e., characteristic length-scales associated with the peaks of structure factors grow initially as $t^{1/3}$. For densities closer to unity, the system exhibits arrested coarsening, with the lack of vacancies and the AA-$B$ repulsion leading to considerable frustration. The system begins to lose ergodicity  [^5] Indeed, for $\rho$ as low as just $0.5$, certain configurations are “frozen”. Examples include the checkerboard pattern (all white squares vacant, alternate black ones filled with $A$ or $B$) and similarly alternate rows or columns. Note that such configurations are also present in the WRLG, as $\rho=0,\pm 1$ appears simply as $n=0,\pm 1$. As a result, the particular density $\rho_{g}$ at which this transition occurs may depend on the initial condition. and exhibits glassy behavior. Given the limitations of our computational resources, systems starting with $\rho_{A,B}\left(\mathbf{s}\right)=\rho/2$ quickly freeze into labyrinthine patterns that resemble those of Turing-Swift-Hohenberg [50]. A rough estimate of this “glass transition” density is $\rho_{g}\sim 0.9$. Typical configurations from these two (expected) phases and the final glassy regime of the $\delta=0$ system are schematically depicted in Fig. 1(b).
@@
The behavior of the driven FLM (at $\delta=1$ here) is more surprising. Though many of its properties are much the same as those in the DWRLG (i.e., disordered at low densities and ordering into regular stripes at high densities), there are some major differences. In particular, it may be argued that there is an extra regime (as $\rho$ is varied) in the FLM, consisting of a phase in-between the analogs of the ordered (high-density) and disordered (low-density) phases of the DWRLG.
@@
At low densities, both the DWRLG and the FLM display a homogeneous, “microemulsion” phase, in which the correlation functions show a periodic structure in the drive direction. For larger $\rho$, both display inhomogeneities. While the DWRLG develops well-defined $A$- or $B$-rich stripes when $\rho>\rho_{c}^{\left(0\right)}\sim 0.6$, stripes begin to emerge in the FLM at a lower value, $\rho_{\ell}\sim 0.4$. In addition, the widths of the stripes in the DWRLG are relatively small and (mostly) independent of $\rho$ for maximal drive ($\delta=1$) [1]. By contrast, the stripes in the FLM that first emerge for $\rho>\rho_{\ell}$ appear to be “irregular” in that their widths fluctuate widely. This regime is apparently not accessible to the DWRLG. Such behavior persists up to an “upper” density, $\rho_{u}\sim 0.65$.
@@
As we increase $\rho$ beyond $\rho_{u}$ and all the way up to the glass transition density $\rho_{g}$, we find properties that more closely resemble those of the DWRLG high-density phase: starting from a uniform initial state, the system quickly develops many stripes with thin, regular widths with a well defined characteristic wavenumber $q^{*}$. We will label this regular stripe regime, $\rho\in\left(\rho_{u},\rho_{g}\right)$. Although investigations with some other initial conditions suggest that there are many long-living metastable states (i.e., not relaxing to regular stripes within the time limits of our runs), a definitive study of them is challenging and beyond the scope of this article [48]. Instead, our main focus here will be the regular and irregular stripe regimes of the FLM. Note that we refrain from referring to these regimes as thermodynamic phases, since extensive simulations with finite-size scaling analysis are necessary to find the critical properties. Deferring such quantitative studies to the future, our goal is to present qualitative observations in the next subsection.
@@
At the highest densities, excluded-volume considerations mark the behavior of the DWRLG and the driven FLM. In the DWRLG, exclusion of AA-$B$ NN pairs prohibits the use of random initial configurations (ICs). In [1], such densities were studied using single-stripe ICs consisting of two blocks (A particles plus vacancies, B particles plus vacancies) separated by empty columns. For the FLM with $\rho>\rho_{g}$, we encounter extremely slow dynamics reminiscent of glass, just as in the $\delta=0$ case. To summarize, these different behaviors and regimes for the driven FLM are depicted in Fig. 1(a).
@@
Figure 1: Summary of steady-states in the FLM. See text for details. Upper panels (a)-(b): Typical charge fluctuation field configurations in a $L=192$ system, also observed in other sizes studied. (a): $\delta=1$. For $\rho<\rho_{\ell}$ (orange background), the system is well mixed and homogeneous, with the structure factors exhibiting a discontinuity at the origin and a peak at $(q_{x},q_{y})=(q^{*},0)$. The intermediate regime (blue background), $\rho_{\ell}<\rho<\rho_{u}$, is marked by irregular stripes which suffer long-wavelength instabilities and display a range of widths. For $\rho>\rho_{u}$ (green background), the stripes are fully developed and long-wavelength instabilities are not sufficient to destroy long-range order in the xx(drive)-direction. This regime is referred to as the regular stripe regime. (b): $\delta=0$. For $\rho<\rho_{c}^{(0)}$ (orange background) the system is well mixed and homogeneous, with the structure factors displaying the Ornstein-Zernike form. For $\rho>\rho_{c}^{(0)}$ (blue background) the system is heterogeneous, exhibiting coarsening dynamics that follows the Lifshitz-Slyozov law. For both $\delta=0$ and $1$, the system undergoes a glassy transition when $\rho>\rho_{g}$ (yellow background). Lower panels (c)-(d): order parameter plots (Eq. (8)) for $\delta=0$ (c) and $\delta=1$ (d); using $L=64$ and averaged over $100$ realizations. The numbered blue marks represent the typical configurations shown in (a) and (b): $\rho=0.2$ (i), $0.38$ (ii), $0.5$ (iii), $0.7$ (iv), $0.9$ (v). Red stars are our crude estimates of the critical density; $\rho_{c}^{(0)}\simeq 0.39$ and $\rho_{\ell}\simeq 0.40$. Red squares are estimates of the glassy transition density, $\rho_{g}\simeq 0.90$. For $\delta=1$, the red triangle marks the transition to regular stripes at the density $\rho_{u}\simeq 0.65$.
@@ II.3 Brief analysis of various regimes
To highlight the novelty of the irregular stripe regime in the driven FLM, we will present a number of approaches to analyzing the $\delta=1$ system. To best identify the various regimes accessible to this model, we focus here on systems with $L=32,64,128,256$ and three densities $\rho=0.39,0.60,0.80$, corresponding to the homogeneous, irregular stripe and regular stripe regimes, respectively. As mentioned above, the typical configurations in the first and last regimes are similar to the “microemulsions” and the striped phases of the DWRLG [2]. By contrast, starting with uniform $\rho_{A,B}$ with overall $\rho\in\left(\rho_{\ell},\rho_{u}\right)$, the system quickly (e.g., $\sim 10^{5}$ MCS for $\rho=0.60$ and $L=256$) forms a relatively large number of thin stripes (of regular width, mainly), as if it is settling into regular stripes. Soon thereafter, these stripes suffer a long-wavelength instability (LWI), which are reminiscent of the Eckhaus instability of convection rolls [51], and break up. They subsequently recombine and, by the end of the runs ($\sim 10^{7}$ MCS), settle into relatively few well-defined (irregular) stripes with a wide range of widths. We illustrate this unusual behavior in Fig. 2 with three snapshots of a $\rho=0.60,\ L=256$ system at about $10^{5},10^{6}$, and $10^{7}$ MCS.
@@
labeled as the density and charge fluctuations: DF $\phi_{+}$ and CF $\phi_{-}$, respectively. Translational invariance of the steady-state allows us to exploit their Fourier transforms (FT)
@@
and study the FT of the correlations, i.e., the structure factors  [^6] Due to the symmetries of the systems we are considering, it can be shown that the cross correlation (associated with $S_{+-}$) vanishes, so that we investigated on these two SFs.
@@
As in the DWRLG, $S_{-}(\mathbf{q})$ is found to peak at a nontrivial wave-vector: $\mathbf{q}^{\ast}\equiv\left(q^{\ast},0\right)$, even in the homogeneous phase, while $S_{+}(\mathbf{q})$ peaks at $2\mathbf{q}^{\ast}$ at high densities (when stripes become more well-established). In fact, the coupling behavior of the DF can be explained in simple terms: since the density field marks the interfaces between A- and B-rich regions, a modulation for the CF must be accompanied by a modulation of twice the frequency in the density. Let us denote the value of $S_{-}$ at the peak as $S_{-}^{\ast}\equiv S_{-}(\mathbf{q}^{\ast})$. Similarly to ordinary equilibrium systems (in $d$ dimensions) with transitions to long-ranged order, we expect this maximum to be $O\left(1\right)$ in the disordered phase, but rise to $O\left(L^{d}\right)$ in the ordered one. Thus, we follow Ref. [1] and define the order parameter as
@@
The two extremes come from a state of complete disorder $\rho_{A,B}\left(\mathbf{s}\right)=\rho/2$, such as those used as initial states in simulations] and one of complete order $\rho_{A,B}\left(\mathbf{s}\right)=0$ or $1$ only]. As the crudest measure of order, it is interesting to report that, e.g., $\Sigma_{-}$ is approximately $0.75\%$ of being maximal for $\rho=0.39$, rising to approximately $15\%$ for the “irregular stripe regime” case of $\rho=0.60$, while reaching about $21\%$ in the high-density $\rho=0.80$ case with regular stripes.
@@
We also demonstrate the different regimes by plotting the $\rho$-dependence of the order parameter, $\Phi\left(\rho\right)$, over a wide range of $\rho$: $\left(0.1,0.9\right)$. In the lower panels of Fig. 1, we illustrate for both $\delta=0$ and $\delta=1$ [Figs. 1(c) and (d), respectively], with $L=64$. In the former, there are clearly two (or three, if the the glassy regime at the highest density is included) regions with very different behavior: $\rho$ below or above $\rho\sim 0.4$. From the associated typical configurations with $L=256$ in (i) and (ii), we identify these regions with the disordered and ordered phases of the $\delta=0$ system, leading to the estimate $\rho_{c}^{\left(0\right)}\simeq 0.39$. By contrast, in Fig. 1(d), there are arguably three (four, including the glassy) regimes for the $\delta=1$ case, from which we estimate $\rho_{\ell}\simeq 0.40$ and $\rho_{u}\simeq 0.65$. Fig. 3(a) further supports these observations by displaying the peak position of the charge fluctuation structure factor, $q^{*}$, as a function of $\rho$ for $\delta=1$ and $L=128$. A clear transition is observed around $\rho\sim 0.65$, where $q^{*}$ exhibits a sharp jump, indicating a qualitative change in behavior within the ordered phase. Figs. 3(b) and (c) show representative steady-state configurations for four selected densities: $\rho=0.475$ (i), $0.625$ (ii), $0.700$ (iii), and $0.800$ (iv), demonstrating that long-range order has developed along the drive direction.
@@
Figure 3: (a): Peak position of the charge fluctuation structure factor (CF SF) along the drive, $q^{*}$, versus the density $\rho$, for $\delta=1$ and $L=128$, averaged over $100$ realizations starting from a uniform initial condition. Color plots: typical field configurations of the CF (b) and the density fluctuation (DF) (c), for: (i) $\rho=0.475$, (ii) $\rho=0.625$, (iii) $\rho=0.7$ and (iv) $\rho=0.8$. Observe the coupling of the DF to the CF, which makes its SF peak position around $2q^{*}$. The color labels for the DF is shifted by $\rho$. Since the DF signal for $\rho=0.475$ is very weak, we mark the interfaces (regions where $\phi_{+}\sim-\rho$) in white.
@@
Another indication of the presence of a new regime between the disordered phase and the regular stripes is that $\Phi\left(\rho,L\right)$ scales with $L$ differently in three regions. As noted above, $\Phi\sim O\left(L^{-d}\right)$ in the disordered phase, rising to $O\left(1\right)$ in the ordered one for standard statistical systems. For example, for a non-conserved Ising model, only in the ordered phase is the magnetization nonzero (in the thermodynamic limit). In the driven FLM, it can be argued that we encounter anomalous scaling in the intermediate, irregular stripe regime. To illustrate this phenomenon, we display plots of $\Phi\left(\rho,L\right)$ for $\delta=0$ [panels (a)] and $\delta=1$ [panels (b)], scaled by various powers of $L$, in Fig. 4. Note that these are semi-log plots, so that data of a wide range of magnitudes can be shown together. We can discern different scaling properties in the three regimes even for a limited range of $L$’s ($32,64,128$). In the lower panels, $\Phi L^{2}$ (i.e., $S_{-}^{\ast}$ itself) is shown and, noting the data collapse in the low density regime $\rho<\rho_{\ell}$ for $\delta=1$ (or $\rho<\rho_{c}^{(0)}$ for $\delta=0$), we may conclude that $\Phi$ has the typical scaling behavior of a disordered phase. In the upper panels, we show that the $\Phi$ values collapse best in the high density regime $\rho>\rho_{u}$ for $\delta=1$ (or $\rho>\rho_{c}^{(0)}$ for $\delta=0$); though considerable finite-size corrections may be at play. By contrast, for $\delta=1$, the middle-right panel, corresponding to intermediate densities $\rho_{\ell}<\rho<\rho_{u}$, displays good data collapse for $\Phi L^{p}$, where $p$ is estimated as $p\sim 0.7$ for the best data collapse. Note that no such collapse occurs for $\delta=0$ (or for any other value of $p$ between 0 and $2$), in the middle-left panel, highlighting the absence of a third regime in the equilibrium FLM. The conclusion is that, for the driven FLM, there is a regime where $\Phi$ scales with a power of $L$ that is “intermediate” between the standard values of 2-2 and 0, corresponding to the usual disordered and ordered phases, respectively. This is our irregular stripe regime.
@@
Figure 4: Data collapse for the order parameter in the different density regimes, when scaled by different powers of $L$. (a): $\delta=0$. (b): $\delta=1$. Dashed lines mark the transition densities. In each case, the different regimes, where data best collapse, are highlighted in red (green) for $\delta=0(1)$. Observe the absence of a third regime in the undriven $\delta=0$ case (middle-(a) panel); see text for details.
@@
Figure 5: The charge fluctuation structure factor, as a function of $q_{x}$ for $q_{y}\,L/2\pi=0,...,10$ (using $L=128$), in the three regimes: microemulsion ($\rho=0.39$) (a), irregular stripes ($\rho=0.60$) (b), and regular stripes ($\rho=0.80$) (c). Note that the $S_{-}$ axis is logarithmic, so that, for the ordered case, there is an extremely sharp “ridge” at $q_{y}=0$. For the irregular stripe regime in (b), this ridge is quite sharp also, but not as extreme.
@@
A further argument for the irregular stripe regime comes from more detailed data involving sections of $S_{-}\left(q_{x},q_{y}\right)$. In Fig. 5(a-c), we present an illustration ($\delta=1,L=128$) of how $S_{-}$ varies with $q_{x}$ for the first decade of $q_{y}\,L/2\pi=0,...,10$ in the three regimes. Deep in the homogeneous (microemulsion) phase, $S_{-}$ roughly follows the Orstein-Zernike form [with two major differences: a discontinuity singularity at $\mathbf{q}\equiv(q_{x},q_{y})=(0,0)$ and a peak at a non-trivial, $L$-independent, $\mathbf{q}^{\ast}=\left(q^{\ast},0\right)$]. In other words, the values of $S_{-}\left(q_{x},q_{y}\right)$ are comparable to $S^{\ast}$ (i.e., the peak is quite broad), as seen in Fig. 5(a). In the opposite extreme, far in the ordered phase, the peak is sharp and narrow, with a value increasing with the system size, e.g., $O\left(L^{2}\right)$ vs. $O\left(1\right)$, in Fig. 5(b). In the irregular stripe regime, the peak value at $\mathbf{q}^{*}=(q^{*},0)$ falls between the two phases and also falls off more slowly as we increase $q_{y}$. To quantitatively appreciate how rapidly $S_{-}$ falls from its maximum values $S_{-}^{\ast}$ as we move away from the ridge, we display the numerical values in Table 1.

### v2

@@ II.2 Results from simulations and comparisons with the lattice gas
To explore the behavior of this model, we carry out simulations on lattices of size $L$ ranging from $32$ to $256$, varying the overall density, $\rho$, from $0.10$ to $0.95$. Here, we report only results from the extreme cases of $\delta=0$ and 11, which exhibit all the key phenomenology of the model at intermediate drive values. The main set of studies employs a uniform initial state with $\rho_{A,B}\left(\mathbf{s}\right)=\rho/2$ and, for measurements in the steady-state, up to $100$ realizations per run, for up to $2\cdot 10^{7}$ MCS. With no drive ($\delta=0$), the behavior is essentially the same as the WRLG, displaying a disordered (phase-separated) state at low (high) densities, with the critical density estimated to be $\rho_{c}^{\left(0\right)}\simeq 0.39$. For $\rho$’s below this value, the structure factors are isotropic and display the standard Ornstein-Zernike form. Near criticality, we expect the properties to be in the Ising class [24] and did not pursue this issue. At densities above $\rho_{c}^{\left(0\right)}$, the usual coarsening dynamics is observed, and we verify [43] that our system follows the Lifshitz-Slyozov law, i.e., characteristic length scales associated with the peaks of structure factors grow initially as $t^{1/3}$. For densities closer to unity, the system exhibits arrested coarsening, with the lack of vacancies and AA-$B$ repulsion leading to considerable frustration. The system begins to lose ergodicity  [^5] Indeed, for $\rho$ as low as just $0.5$, certain configurations are “frozen”. Examples include the checkerboard pattern (all white squares vacant, alternate black ones filled with $A$ or $B$) and similarly alternate rows or columns. Note that such configurations are also present in the WRLG, as $\rho=0,\pm 1$ appears simply as $n=0,\pm 1$. As a result, the particular density $\rho_{g}$ at which this transition occurs may depend on the initial condition. and exhibits glassy behavior. Systems starting with $\rho_{A,B}\left(\mathbf{s}\right)=\rho/2$ quickly freeze into labyrinthine patterns that resemble those of the Turing-Swift-Hohenberg model [45]. A rough estimate of the “glass transition” density is $\rho_{g}\sim 0.9$. Typical configurations from these two (expected) phases and the final glassy regime of the $\delta=0$ system are schematically depicted in Fig. 1(b).
@@
The behavior of the driven FLM (at $\delta=1$ here) is more surprising. Though many of its properties are much the same as those in the DWRLG (i.e., disordered at low densities and ordering into regular stripes at high densities), there are some major differences. In particular, it may be argued that there is an extra regime (as $\rho$ is varied) in the FLM, consisting of a phase in between the analogs of the ordered (high-density) and disordered (low-density) phases of the DWRLG.
@@
At low densities, both the DWRLG and the FLM display a homogeneous, “microemulsion” phase, in which the correlation functions show a periodic structure in the drive direction. For larger $\rho$, both display inhomogeneities. While the DWRLG develops well defined $A$- and $B$-rich stripes when $\rho>\rho_{c}^{\left(0\right)}\sim 0.6$, stripes begin to emerge in the FLM at a lower value, $\rho_{\ell}\sim 0.4$. In addition, the widths of the stripes in the DWRLG are relatively small and (mostly) independent of $\rho$ for maximal drive ($\delta=1$) [1]. By contrast, the stripes in the FLM that first emerge for $\rho>\rho_{\ell}$ appear to be “irregular” in that their widths fluctuate widely. This regime is apparently not accessible to the DWRLG. Such behavior persists up to an “upper” density, $\rho_{u}\sim 0.65$.
@@
As we increase $\rho$ beyond $\rho_{u}$ and all the way up to the glass transition density $\rho_{g}$, we find properties that more closely resemble those of the DWRLG high-density phase: starting from a uniform initial state, the system quickly develops many stripes with thin, regular widths with a well defined characteristic wavenumber $q^{*}$. We will label this the regular stripe regime, $\rho\in\left(\rho_{u},\rho_{g}\right)$. Although investigations with some other initial conditions suggest that there are many long-lived metastable states (i.e., not relaxing to regular stripes within the time limits of our runs), a definitive study of them is challenging and beyond the scope of this article [43]. Instead, our main focus here will be the regular and irregular stripe regimes of the FLM. Although the various regimes encountered in our simulations and numerical integration appear to represent distinct phases, we caution the reader that the term is used somewhat informally. Deferring more systematic studies (including finite-size scaling analyses) required to establish the existence of well defined phases (and associated phase transitions) to future work, our goal is to present qualitative observations in the next subsection.
@@
At the highest densities, excluded-volume considerations mark the behavior of the DWRLG and the driven FLM. In the DWRLG, exclusion of AA-$B$ NN pairs prohibits the use of random initial configurations. In [1], such densities were studied using single-stripe initial configurations consisting of two blocks (A particles plus vacancies, B particles plus vacancies) separated by empty columns. For the FLM with $\rho>\rho_{g}$, we encounter extremely slow dynamics reminiscent of glass, just as in the $\delta=0$ case. Comparing panels (c) and (d) of Fig. 1, we note that the precipitous drop in $\Phi$, marking the onset of the glassy regime, occurs at nearly the same density in the undriven and maximally-driven cases, suggesting that this kinetic phenomenon is independent of the drive, and occurs well before ordering can develop. To summarize, these different behaviors and regimes for the driven FLM are depicted in Fig. 1(a).
@@
Figure 1: Summary of steady-states in the FLM. See text for details. Upper panels (a)-(b): Typical charge fluctuation field configurations in a $L=192$ system, also observed in the other sizes studied. (a): $\delta=1$. For $\rho<\rho_{\ell}$ (orange background), the system is well mixed and homogeneous, with the structure factors exhibiting a discontinuity at the origin and a peak at $(q_{x},q_{y})=(q^{*},0)$. The intermediate regime (blue background), $\rho_{\ell}<\rho<\rho_{u}$, is marked by irregular stripes which suffer long-wavelength instabilities and display a range of widths. For $\rho>\rho_{u}$ (green background), the stripes are fully developed and long-wavelength instabilities are not sufficient to destroy long-range order in the xx(drive)-direction. This regime is referred to as the regular stripe regime. (b): $\delta=0$. For $\rho<\rho_{c}^{(0)}$ (orange background) the system is well mixed and homogeneous, with the structure factors displaying the Ornstein-Zernike form. For $\rho>\rho_{c}^{(0)}$ (blue background) the system is heterogeneous, exhibiting coarsening dynamics that follows the Lifshitz-Slyozov law. For both $\delta=0$ and $1$, the system undergoes a glassy transition when $\rho>\rho_{g}$ (yellow background). Lower panels (c)-(d): order parameter plots [Eq. (8)] for $\delta=0$ (c) and $\delta=1$ (d); using $L=64$ and averaged over $100$ realizations. The blue Roman numerals (i-v) near $\rho=0.2,0.4,0.5,0.7,0.9$ correspond to the typical configurations shown in (a) and (b). Red stars are our crude estimates of the critical density; $\rho_{c}^{(0)}\simeq 0.39$ and $\rho_{\ell}\simeq 0.40$. Red squares are estimates of the glassy transition density, $\rho_{g}\simeq 0.90$. For $\delta=1$, the red triangle marks the transition to regular stripes at the density $\rho_{u}\simeq 0.65$. A movie related to this figure is available in [46].
@@ II.3 Brief analysis of various regimes
To highlight the novelty of the irregular stripe regime in the driven FLM, we will present a number of approaches to analyzing the $\delta=1$ system. To best identify the various regimes accessible to this model, we focus here on systems with $L=32,64,128,256$ and three densities, $\rho=0.39,0.60,0.80$, corresponding to the homogeneous, irregular stripe and regular stripe regimes, respectively. As mentioned above, the typical configurations in the first and last regimes are similar to the “microemulsions” and the striped phases of the DWRLG [2]. By contrast, starting with uniform $\rho_{A,B}$ with overall $\rho\in\left(\rho_{\ell},\rho_{u}\right)$, the system quickly (e.g., $\sim 10^{5}$ MCS for $\rho=0.60$ and $L=256$) forms a relatively large number of thin stripes (of regular width, mainly), as if it were settling into regular stripes. Soon thereafter, these stripes suffer a long-wavelength instability, reminiscent of the Eckhaus instability of convection rolls [47], and break up. They subsequently recombine and, by the end of the runs ($\sim 10^{7}$ MCS), settle into relatively few well-defined (irregular) stripes with a wide range of widths. We illustrate this unusual behavior in Fig. 2 with three snapshots of a $\rho=0.60,\ L=256$ system at about $10^{5},10^{6}$, and $10^{7}$ MCS.
@@
labeled as the density (DF) and charge (CF) fluctuations: $\phi_{+}$ and $\phi_{-}$, respectively. Translational invariance of the steady-state allows us to exploit their Fourier transforms
@@
and study the transform of the correlations, i.e., the structure factors (SF)  [^6] Due to the symmetries of the systems we are considering, it can be shown that the cross correlation (associated with $S_{+-}$) vanishes, so that we investigated on these two SFs.
@@
As in the DWRLG, $S_{-}(\mathbf{q})$ is found to peak at a nonzero wave-vector: $\mathbf{q}^{\ast}\equiv\left(q^{\ast},0\right)$, even in the homogeneous phase, while $S_{+}(\mathbf{q})$ peaks at $2\mathbf{q}^{\ast}$ at high densities (when stripes become more well-established). In fact, the doubling of the frequency for the DF can be explained in simple terms: since the density field marks the interfaces between A- and B-rich regions, a modulation for the CF must be accompanied by a modulation of twice the frequency in the density. Let us denote the value of $S_{-}$ at the peak as $S_{-}^{\ast}\equiv S_{-}(\mathbf{q}^{\ast})$. As in ordinary equilibrium systems (in $d$ dimensions) with transitions to long-range order, we expect this maximum to be $O\left(1\right)$ in the disordered phase, but rise to $O\left(L^{d}\right)$ in the ordered one. Thus, we follow Ref. [1] and define the order parameter as
@@
The two extremes come from a state of complete disorder $\rho_{A,B}\left(\mathbf{s}\right)=\rho/2$, such as those used as initial states in simulations] and one of complete order $\rho_{A,B}\left(\mathbf{s}\right)=0$ or $1$ only]. As the crudest measure of order, it is interesting to report that, e.g., $\Sigma_{-}$ is approximately $0.75\%$ of being maximal for $\rho=0.39$, rising to approximately $15\%$ for the irregular stripe regime case of $\rho=0.60$, while reaching about $21\%$ in the high-density $\rho=0.80$ case with regular stripes.
@@
We also demonstrate the different regimes by plotting the $\rho$-dependence of the order parameter, $\Phi\left(\rho\right)$, over a wide range of $\rho$: $\left(0.1,0.9\right)$. In the lower panels of Fig. 1, we illustrate for both $\delta=0$ and $\delta=1$ [Figs. 1(c) and (d), respectively], with $L=64$. In the former, there are clearly two (or three, if the the glassy regime at the highest density is included) regions with very different behavior: $\rho$ below or above $\rho\sim 0.4$. From the associated typical configurations with $L=256$ in (i) and (ii), we identify these regions with the disordered and ordered phases of the $\delta=0$ system, leading to the estimate $\rho_{c}^{\left(0\right)}\simeq 0.39$. By contrast, in Fig. 1(d), there are arguably three (four, including the glassy) regimes for the $\delta=1$ case, from which we estimate $\rho_{\ell}\simeq 0.40$ and $\rho_{u}\simeq 0.65$. We note that the order parameter in the ordered regimes of the driven ($\delta=1$) system is over an order of magnitude larger than in the corresponding regimes of the undriven case. Fig. 3(a) further supports these observations by displaying the peak position of the charge fluctuation structure factor, $q^{*}$, as a function of $\rho$ for $\delta=1$ and $L=128$. A clear transition is observed around $\rho\sim 0.65$, where $q^{*}$ exhibits a sharp jump, indicating a qualitative change in behavior within the ordered phase. Figs. 3(b) and (c) show representative steady-state configurations for four selected densities: $\rho=0.475$ (i), $0.625$ (ii), $0.700$ (iii), and $0.800$ (iv), demonstrating that long-range order has developed along the drive direction.
@@
Figure 3: (a): Peak position of the charge fluctuation structure factor (CF SF) along the drive, $q^{*}$, versus the density $\rho$, for $\delta=1$ and $L=128$, averaged over $100$ realizations starting from a uniform initial condition. Color plots: typical field configurations of the CF (b) and the density fluctuation (DF) (c), for: (i) $\rho=0.475$, (ii) $\rho=0.625$, (iii) $\rho=0.7$ and (iv) $\rho=0.8$. Observe that the DF has minima at the interfaces of the CF. The color map scale for the DF is shifted by $\rho$. Since the DF signal for $\rho=0.475$ is very weak, we mark the interfaces (regions where $\phi_{+}\sim-\rho$) in white.
@@
Another indication of the presence of a new regime between the disordered and regular-stripe regimes is that $\Phi\left(\rho,L\right)$ scales differently with $L$ in three regions. As noted above, $\Phi\sim O\left(L^{-d}\right)$ in the disordered phase, rising to $O\left(1\right)$ in the ordered one for standard statistical systems. For example, for a non-conserved Ising model, the magnetization is nonzero only in the ordered phase (in the thermodynamic limit). In the driven FLM, it can be argued that we encounter anomalous scaling in the intermediate, irregular stripe regime. To illustrate this phenomenon, we display plots of $\Phi\left(\rho,L\right)$ for $\delta=0$ [panels (a)] and $\delta=1$ [panels (b)], scaled by various powers of $L$, in Fig. 4. Note that these are semi-log plots, so that data of a wide range of magnitudes can be shown together. We can discern different scaling properties in the three regimes even for a limited range of $L$’s ($32,64,128$). In the lower panels, $\Phi L^{2}$ (i.e., $S_{-}^{\ast}$ itself) is shown and, noting the data collapse in the low density regime, $\rho<\rho_{\ell}$, for $\delta=1$ (or $\rho<\rho_{c}^{(0)}$ for $\delta=0$), we may conclude that $\Phi$ has the typical scaling behavior of a disordered phase. In the upper panels, we show that the $\Phi$ values collapse best in the high density regime $\rho>\rho_{u}$ for $\delta=1$ (or $\rho>\rho_{c}^{(0)}$ for $\delta=0$), though considerable finite-size corrections may be at play. By contrast, for $\delta=1$, the middle-right panel, corresponding to intermediate densities $\rho_{\ell}<\rho<\rho_{u}$, displays good data collapse for $\Phi L^{p}$, where $p$ is estimated as $p\sim 0.7$ for the best data collapse. Note that no such collapse occurs for $\delta=0$ (or for any other value of $p$ between 0 and $2$), in the middle-left panel, highlighting the absence of a third regime in the equilibrium FLM. The conclusion is that, for the driven FLM, there is a regime where $\Phi$ scales with a power of $L$ that is “intermediate” between the standard values of 2-2 and 0, corresponding to the usual disordered and ordered phases, respectively. This is our irregular stripe regime.
@@
Figure 4: Data collapse for the order parameter in the different density regimes, when scaled by different powers of $L$. (a): $\delta=0$. (b): $\delta=1$. Dashed lines mark the transition densities. In each case, the different regimes, where data best collapse, are highlighted in red (green) for $\delta=0(1)$. Error bars are mainly smaller than the size of the symbols. Observe the absence of a third regime in the undriven $\delta=0$ case (middle-(a) panel); see text for details.
@@
Figure 5: $\delta=1$: The charge fluctuation structure factor, as a function of $q_{x}$ for $q_{y}\,L/2\pi=0,...,10$ (using $L=128$), in the three regimes: microemulsion ($\rho=0.39$) (a), irregular stripes ($\rho=0.60$) (b), and regular stripes ($\rho=0.80$) (c). Note that the $S_{-}$ axis is logarithmic, so that, for the ordered case, there is an extremely sharp “ridge” at $q_{y}=0$. For the irregular stripe regime in (b), this ridge is quite sharp also, but not as extreme.
@@
A further argument for the irregular stripe regime comes from more detailed data involving sections of $S_{-}\left(q_{x},q_{y}\right)$. In Fig. 5(a-c), we present an illustration ($\delta=1,L=128$) of how $S_{-}$ varies with $q_{x}$ for the first decade of $q_{y}\,L/2\pi=0,...,10$ in the three regimes. Deep in the homogeneous (microemulsion) regimes, $S_{-}$ roughly follows the Orstein-Zernike form [with two major differences: a discontinuity singularity at $\mathbf{q}\equiv(q_{x},q_{y})=(0,0)$ and a peak at a nonzero, $L$-independent, $\mathbf{q}^{\ast}=\left(q^{\ast},0\right)$]. In other words, the values of $S_{-}\left(q_{x},q_{y}\right)$ are comparable to $S^{\ast}$ (i.e., the peak is quite broad), as seen in Fig. 5(a). In the opposite extreme, deep in the ordered regime, the peak is sharp and narrow (with a value increasing with the system size), as seen in Fig. 5(c). In the irregular stripe regime [Fig. 5(b)], the peak value at $\mathbf{q}^{*}=(q^{*},0)$ falls between the two phases and also falls off more slowly as we increase $q_{y}$.cTo quantitatively appreciate how rapidly $S_{-}$ falls from its maximum values $S_{-}^{\ast}$ as we move away from the ridge, we display the numerical values in Table 1.